OpenStudy (anonymous):

-8x-(3-x)=-15 solve for x thanks

OpenStudy (akshay_budhkar):

you want the steps?

OpenStudy (anonymous):

can you solve step by step thanks

OpenStudy (mimi_x3):

-8x-3+x = -15 -7x-3=-15 -7x = -12 x = 7/12 (not sure if im right though)

OpenStudy (akshay_budhkar):

-8x-3+x=-15 -7x=-12 x=12/7

OpenStudy (mimi_x3):

12/7*

OpenStudy (akshay_budhkar):

its 12/7

OpenStudy (anonymous):

oops omg i copied the problem wrong lol its -8x-8(3-x)=-15 sorry lol

OpenStudy (mimi_x3):

lols i'll do it again

OpenStudy (anonymous):

thanks lol

OpenStudy (akshay_budhkar):

no solution

OpenStudy (akshay_budhkar):

it cant be.. x vanishes

OpenStudy (anonymous):

thanks i got the same thing -24=-15

OpenStudy (mimi_x3):

yeah bcos x is zero

OpenStudy (anonymous):

thanks guys

OpenStudy (mimi_x3):

yw xDD

OpenStudy (mimi_x3):

post another question im so~ bored xDD

OpenStudy (akshay_budhkar):

lol mimi try 2-2

OpenStudy (mimi_x3):

LOLS 0 common sense

OpenStudy (anonymous):

lol you guys are smart at algebra 2

OpenStudy (akshay_budhkar):

i am smart at maths i believve. you can bank on me! lol

OpenStudy (mimi_x3):

nah im dumb lmao

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

lol psh yeah right..

OpenStudy (mimi_x3):

lols , im srs the questions yu posted were easy manss

OpenStudy (anonymous):

lol so ur calling me dumb Dx

OpenStudy (mimi_x3):

lolss , woops nah

OpenStudy (anonymous):

lol im kidding ;p

OpenStudy (anonymous):

no one here is dumb we're learning ;p

OpenStudy (mimi_x3):

kk (:

OpenStudy (mimi_x3):

yeah LOL

OpenStudy (anonymous):

ill post another question soon maybe in like 5 mins doing my summer packet lol

OpenStudy (mimi_x3):

kk lmao